Find what the line xy-x-y+1=0 will become when the origin is shifted to the point (1,1). 

Grade:11

1 Answers

Susmita
5 years ago
You can just replace x by x-1 and y by y-1.
xy-x-y+1=0
Or,(x-1)(y-1)-(x-1)-(y-1)+1=0
Or,xy-x-y+1-x+1-y+1+1=0
Or,xy-2x-2y+3=0